# Contrastive learning of T cell receptor representations

### v1

@@ Abstract
Antigen-specific T cells play important protective and pathogenic roles in human disease [1]. The recognition of peptides presented on major histocompatibility complexes (pMHCs) by T cell receptors (TCRs) determines the specificity of cellular immune responses [2]. Hyperdiverse TCRs are generated during T cell development in the thymus by genetic recombination of germline-encoded V, D (for TCR) and J gene segments with additional diversification through insertions and deletions of non-template nucleotides at gene segment junctions.
@@
The most direct approach for applying machine learning to TCR specificity prediction has been to train pMHC-specific models that take an arbitrary TCR and predict binding [7, 8, 9, 10, 11, 12]. More ambitiously, model architectures have been proposed that can in principle generalise predictions to arbitrary pMHCs as well [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Independent benchmarking studies have shown that both approaches are effective for predicting TCR binders against pMHCs for which many TCRs have been experimentally determined [25], but generalisation to pMHCs not seen during training has largely remained elusive [26] and prediction accuracy is limited for pMHCs with few known binders [27]. This severely limits the utility of current predictive tools given that only of the possible pMHCs are currently annotated with any TCRs in VDJdb [28], and given that for of them less than 100 specific TCRs are known.
@@ I.1 Benchmarking PLM embeddings on TCR specificity prediction
To our surprise, we found that TCR-BERT, ESM2, and ProtBert all fail to outperform the baseline sequence alignment method (CDR3 Levenshtein) and are significantly inferior to TCRdist (Figs. 1d / S1). A repeat of the benchmarking with a broader set of epitopes obtained by including post-processed 10x Genomics whitepaper data [42], recapitulated these results, demonstrating the robustness of our findings (Fig. S2). In contrast to existing PLMs, SCEPTR performs on par with or better than TCRdist. For a reference set of size 200, SCEPTR performs best among all models for five out of six tested peptides (Table SI).
We additionally compared models using the average distance between a query TCR and all references, instead of only the nearest neighbour (Fig. S3). In this case, SCEPTR outperforms other models by an even wider margin. Interestingly, all models perform worse compared to their nearest neighbour counterpart. This finding might be explained mechanistically by the multiplicity of viable binding solutions with distinct sequence-level features, which are thought to make up pMHC-specific TCR repertoires [44, 45].
@@ I.2 Autocontrastive learning as a pre-training strategy
There are several well-known variants of this learning approach. In supervised contrastive learning, positive pairs are generated by sampling observations known to belong to the same class (Fig. 2d top). In the context of TCRs, we can define positive pairs to be TCRs annotated to interact with the same pMHC, in which case contrastive learning regresses distances between TCR pairs to their probabilities of co-specificity. Autocontrastive learning approximates such positive pairs through data augmentation by generating two independent views of the same observation by passing it twice through the model (Fig. 2d bottom).
Given the scarcity of available labelled data, we opted to use the autocontrastive approach for purely unsupervised PLM pre-training (see Sec. I.5 for an application of supervised contrastive learning, more similar to other recent applications of contrastive learning to TCRs [49, 50]). We generate different “views” of a TCR by dropout noise as is standard in NLP [51], but additionally adopted a censoring strategy inspired by masked-language modeling that randomly removes a proportion of residues or even complete or chains. In contrast to the only other study known to us having explored the application of autocontrastive learning to TCRs [52], we trained SCEPTR on all six hypervariable loops of the full paired chain TCR, as all contribute to TCR-pMHC specificity [45]. That being said, our chain dropping procedure during censoring ensures that single chain data are also in distribution for the model, giving SCEPTR flexibility for downstream applications with bulk sequenced TCR repertoires.
@@ I.3 Ablation studies
The MLM-only variant underperforms compared to both SCEPTR and TCRdist, demonstrating that autocontrastive learning is a necessary ingredient for the increased performance of SCEPTR in few-shot specificity prediction (Figs. 3a / S4a).
@@
While SCEPTR’s <cls> embeddings achieve the best results, the autocontrastive average-pooling variant still performs on par with TCRdist (Figs. 3b / S4b).
@@
This variant is trained on a size-matched set unlabelled TCRs generated by OLGA [54], a probabilistic model of VDJ recombination.
@@
This variant is trained on the same set of TCRs as the original model, but the / chain pairing is randomised.
We find that pairing information and the use of real post-selection repertoires slightly improves SCEPTR’s performance, but variants trained on alternative datasets still match TCRdist performance (Figs. 3c / S4c). The synthetic data models the statistics of VDJ recombination and therefore estimates a pre-thymic TCR distribution. As such, SCEPTR’s slight performance improvement using peripheral blood TCRs hints at a potential ability of language modelling to extract information from imprints left by antigen-driven selection in memory repertoires [44]. Similarly, the slight performance increase on paired TCR data suggests some ability to extract information from pairing biases [55].
@@
The results demonstrate that taking into account all CDR loops leads to a performance gain as expected (Figs. 3d / S4d). We also see that autocontrastive learning even when restricted to CDR3s leads to a substantial performance gain, helping the autocontrastive CDR3 variant achieve similar performance to the full-input MLM-only variant (Fig. 3a,d).
I.4 Learning features within embedding spaces
So far we have focused on nearest-neighbor prediction using PLM embeddings as the most direct test of data-efficient transfer learning that works with as little as a single reference sequence. If slightly more data is available, another approach is to train simple supervised predictors atop PLM embeddings. To test how much such training can improve prediction performance, we trained linear support vector classifiers (SVC) on the PLM embeddings provided by different models. In each instance, we trained the classifier to distinguish reference TCRs from 1000 randomly sampled background TCRs (see methods III.4).
We find that the SVC predictors for ProtBert, ESM2 and TCR-BERT all perform better than their nearest-neighbour counterparts, but still worse than SCEPTR’s nearest neighbour predictions (Fig. 4 / S6). We also trained an SVC atop SCEPTR, which did not lead to further improvement upon the nearest-neighbor prediction (Fig. S6). These findings highlight how in the low data regime typical of most pMHCs, misalignment of pre-training to downstream tasks can only be partially remediated by training on reference TCRs.
@@ I.5 Supervised contrastive learning as a fine-tuning strategy
For this task we took all the TCRs annotated against the target pMHCs from our labelled TCR dataset, and split them into a training, a validation, and a testing set. We ensured that no study used for training or validation contributed any data to the test set, so that the fine-tuned model would not be able to achieve good performance simply by exploiting inter-dataset biases. The training set included 200 binders against each target pMHC, totalling to 1200 TCRs. The rest of the TCRs from the same studies were used to construct the validation set. TCRs from all remaining studies were used for the testing set, which comprised of 5670 TCRs. SCEPTR was fine-tuned on the training set with supervised contrastive learning, using the validation loss for early stopping (appendix III.5).
We used the framework from section I.1 to benchmark the performance of fine-tuned SCEPTR, using the training set as the references. The results show that fine-tuning can greatly improve the ability of the model to discriminate between specificities (Fig. 5). Improvements are most noticeable for the pMHCs against which other methods achieve relatively low performance. However, model performance degrades with respect to unseen pMHCs (Fig. S5), perhaps unsurprisingly given the very limited number of pMHCs represented in the training data. Interestingly, unlike other models (Fig. S3), fine-tuned SCEPTR makes better inferences by measuring the average distance between a query TCR and all reference TCRs instead only the nearest TCR (Fig. S7). This suggests an ability of supervised contrastive fine-tuning to help the model discover the commonalities between the multiple different binding solutions thought to exist for each pMHC.
@@ II Discussion
A limitation of our study is that we did not undertake a complete exploration of training and architectural hyperparameters. We envisage multiple avenues that may improve SCEPTR further. Firstly, training could be made more efficient by optimising the distribution of masked/dropped tokens during pre-training, taking into account the variable relevance of different parts of the sequence in determining specificity [45]. Secondly, as certain sequence motifs appear recurrently (e.g. CDR3 loops often begin with CAS), a more intelligent tokenisation scheme could offload learning of these primary sequence statistics into the tokenisation process.
@@
Uniformity incentivises the model to make use of the full representation space, while alignment minimises the expected distance between positive pairs (e.g. co-specific TCRs) [47]. From this view, contrastive learning on adaptive immune receptor data encourages PLMs to undo the large-scale distributional biases created by VDJ recombination through the uniformity term, while helping to identify features relating to TCR (co-)specificity via the alignment term. While autocontrastive learning approximates the alignment term through the generation of pairs of views, it still provides a direct empirical estimate for the uniformity term. Thus, a key benefit of autocontrastive learning may be that it reduces the confounding effects of VDJ recombination in embedding space.
@@
Another interesting avenue is the use of labels other than pMHC specificity – such as phenotypic annotations from single-cell data – as additional supervised contrastive training signals. Finally, while supervised contrastive learning does not currently lead to generalisable learning beyond training pMHCs, we expect a transition towards generalisation as larger volumes of specificity-labelled TCR data become available, as has been the case with supervised contrastive learning in other fields [64, 65, 66, 67, 51]. Finally, while the focus of this work given current data limitations has been on learning TCR embeddings, contrastive learning may also help us learn effective joint TCR-pMHC embeddings in the future when the joint space (and particularly the pMHC space) is better sampled, and thus ultimately enable the zero-shot prediction of TCR-pMHC specificity.
@@ III.1 Model benchmarking
For each pMHC, we varied the number of reference TCRs where . Within each model-pMHC--shot combination, we benchmarked multiple reference-test splits of the data to ensure robustness. For , we benchmarked every possible split. For , we benchmarked 100 random splits, where we ensured that the same splits were used across all models to reduce extraneous variance. Variations in model performance across different splits were used to generate confidence intervals for differences in performance, for example in figure S1.
@@ III.2 SCEPTR architecture
SCEPTR uses a simple, non-trainable embedder module, where a one-hot vector is used to encode token identity (22 dimensions for 20 amino acids plus special tokens <cls> and <mask>), and token positions are specified by first one-hot encoding the containing CDR loop number (6 dimensions), then encoding the token’s relative position within the loop as a single scalar variable (Fig. 2b). This results in initial token embeddings in , which are passed through a trainable linear projection onto . SCEPTR’s self-attention stack then operates at this fixed dimensionality (Fig. S11). SCEPTR’s self-attention stack comprises three layers, each with eight attention heads and a feed-forward dimensionality of 256, and is thus substantially simpler than existing models. Our tests suggest that relative position embedding helps SCEPTR learn better calibrated TCR co-specificity rules (see appendix B).
@@ III.3.1 Data
The unlabelled paired-chain TCR sequences used to pre-train SCEPT were taken from a study by Tanno et al. [68], which provides 965,523 unique clonotypes sampled from the blood of 15 healthy human subjects. As opposed to traditional single-cell sequencing, Tanno et al. used a ligation-based sequencing methods to resolve which chains paired with which chains. To mitigate potential noise from incorrect chain pairing, we applied an extra processing step to remove clonotypes that shared the same nucleotide sequence for either the or the chain, as previously described [44]. After filtering for functional TCRs using tidytcells, a TCR gene symbol standardiser [69], we retained 842,683 distinct clonotypes.

### v2

@@ Abstract
Antigen-specific T cells play important protective and pathogenic roles in human disease [1]. The recognition of peptides presented on major histocompatibility complexes (pMHCs) by T cell receptors (TCRs) determines the specificity of cellular immune responses [2]. Hyperdiverse TCRs are generated during T cell development in the thymus by genetic recombination of germline-encoded V, D (for TCR) and J gene segments with additional diversification by trimming of the germline and insertions of non-template nucleotides at gene segment junctions.
@@
The most direct approach for applying machine learning to TCR specificity prediction has been to train pMHC-specific models that take an arbitrary TCR and predict binding [7, 8, 9, 10, 11, 12]. More ambitiously, model architectures have been proposed that can in principle generalise predictions to arbitrary pMHCs as well [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Independent benchmarking studies have shown that both approaches are effective for predicting TCR binders against pMHCs for which many TCRs have been experimentally determined [25], but generalisation to pMHCs not seen during training has largely remained elusive [26] and prediction accuracy is limited for pMHCs with few known binders [27]. This severely limits the utility of current predictive tools given that to date, only of the possible pMHCs are annotated with any TCRs in VDJdb [28], and given that for of them less than 100 specific TCRs are known.
@@ I.1 Benchmarking PLM embeddings on TCR specificity prediction
To our surprise, we found that TCR-BERT, ESM2, and ProtBert all fail to outperform the baseline sequence alignment method (CDR3 Levenshtein) and are significantly inferior to TCRdist (Figs. 1d / S1). A repeat of the benchmarking with a broader set of epitopes obtained by including post-processed 10x Genomics whitepaper data [42] recapitulated these results, demonstrating the robustness of our findings (Fig. S2). In contrast to existing PLMs, SCEPTR performs on par with or better than TCRdist (Figs. 1d / S1, Table SI). For a reference set of size 200, SCEPTR performs better than TCRdist for five out of six tested peptides (, binomial test), and for all six peptides when compared to all other models (, binomial test).
We additionally compared models using the average distance between a query TCR and all references, instead of only the nearest neighbour (Fig. S3). In this case, SCEPTR outperforms other models by an even wider margin. Interestingly, all models perform worse compared to their nearest neighbour counterpart. This finding might be explained mechanistically by the multiplicity of viable binding solutions with distinct sequence-level features, which are thought to make up pMHC-specific TCR repertoires [44, 45]. We hypothesized that the poor performance of prior PLMs might be overcome by learning projections from their high-dimensional representation space that align better with the TCR co-specificity prediction task. We thus used the embeddings as input to linear support vector classifiers (Appendix B). Optimised linear probing does improve prediction performance (Fig. S4), but they remain inferior to nearest neighbour predictions using SCEPTR further illustrating the usefulness of SCEPTR embeddings for data-efficient transfer learning.
In some use cases, we may want to apply SCEPTR to the analysis of single chain TCR data. In benchmarking on either or chain reference data alone prediction accuracy drops somewhat regardless of distance measure (Fig. S5), as expected given that both receptor chains provide non-redundant information about specificity [45]. Importantly though, SCEPTR distances also provide comparable or better prediction accuracy than TCRdist on the single chain level.
@@ I.2 Autocontrastive learning as a pre-training strategy
There are several well-known variants of this learning approach. In supervised contrastive learning, positive pairs are generated by sampling observations known to belong to the same class (Fig. 2d top). In the context of TCRs, we can define positive pairs to be TCRs annotated to interact with the same pMHC, in which case contrastive learning regresses distances between TCR pairs to their probabilities of co-specificity. Autocontrastive learning approximates such positive pairs through data augmentation by generating two independent “views”’ of the same observation (Fig. 2d bottom).
Given the scarcity of available labelled data, we opted to use the autocontrastive approach for purely unsupervised PLM pre-training (see Sec. I.5 for an application of supervised contrastive learning, more similar to other recent applications of contrastive learning to TCRs [49, 50]). For this pre-training, we used data on close to a million unique paired TCRs obtained by Tanno et al. [51], which represents one the largest collections of TCRs from a single study collected to date. This data was filtered and standardized as described in methods III.3. We generate different “views” of a TCR by dropout noise as is standard in NLP [52], but additionally adopted a censoring strategy inspired by masked-language modeling that randomly removes a proportion of residues or even complete or chains. In contrast to the only other study known to us having explored the application of autocontrastive learning to TCRs [53], we trained SCEPTR on all six hypervariable loops of the full paired chain TCR, as all contribute to TCR-pMHC specificity [45]. That being said, our chain dropping procedure during censoring ensures that single chain data are also in distribution for the model, giving SCEPTR flexibility for downstream applications with bulk sequenced TCR repertoires (Fig. S5).
@@ I.3 Ablation studies
The MLM-only variant underperforms compared to both SCEPTR and TCRdist, demonstrating that autocontrastive learning is a necessary ingredient for the increased performance of SCEPTR in few-shot specificity prediction (Fig. 3a).
@@
While SCEPTR’s <cls> embeddings achieve the best results, the autocontrastive average-pooling variant still performs on par with TCRdist (Fig. 3b).
@@
This variant is trained on a size-matched set unlabelled TCRs generated by OLGA [55], a probabilistic model of VDJ recombination. This synthetic data models only the recombination statistics and thus estimates the TCR distribution without taking into account the imprints of thymic and peripheral selection found in real repertoires [44].
@@
This variant is trained on the same set of TCRs as the original model, but the / chain pairing is randomised, thus removing pairing biases [56].
We find that SCEPTR trained on synthetic or shuffled data performs worse for five out of six pMHCs (, binomial test), but differences in AUROCs between model variants are small and regardless of training data SCEPTR performs on par with TCRdist (Fig. 3c).
@@
The results demonstrate that taking into account all CDR loops leads to a performance gain as expected (Fig. 3d). We also see that autocontrastive learning even when restricted to CDR3s leads to a substantial performance gain, helping the autocontrastive CDR3 variant achieve similar performance to the full-input MLM-only variant.
I.4 Comparison of SCEPTR embeddings to alignment-based TCR similarity
To gain insights into what SCEPTR has learned during pre-training, we compared its embedding distances to alignment-based distances as calculated by TCRdist. To do so, we calculated all pairwise distances between one thousand TCRs randomly sampled from the testing partition of the Tanno et al. dataset [51] (held out during SCEPTR pre-training) using both models. We find that SCEPTR and TCRdist distances are clearly correlated (Fig. 4a). This shows that in parts the success of SCEPTR can be understood by its embedding distances providing good alignment-free approximations to traditional sequence similarity measures. Yet, there is also substantial variability between both measures for many pairs, and an inspection of such discordant pairs can provide insights into how metrics differ.
First, we noticed that pairs of sequences judged to be similar by SCEPTR but not TCRdist were less likely to be generated during VDJ recombination (Fig. 4b). We found that among sequence pairs judged to be similar by SCEPTR, TCRdist distance was strongly negatively correlated with (Figs. 4c / S6). This implies that SCEPTR embeds TCRs closer to each other if they are in regions of sequence space that are less densely sampled by the generative distribution. As argued in detail in the Discussion, this property of SCEPTR embeddings is expected on theoretical grounds due to the loss function used for contrastive learning. This property might enable SCEPTR embeddings to capture the intuition that finding close-by nearest neighbours is more surprising for sequences with low and thus more informative compared to similarity between highly probable TCRs.
Second, we noticed that a high similarity on a single chain tended to be sufficient for a small SCEPTR distance (Fig. S7). To quantify this effect, we analysed how SCEPTR distances correlate with different ways of averaging the and chain TCRdist distances into a paired chain measure. We found that SCEPTR distances correlate more closely with the minimum distance of the two chains rather than their arithmetic mean (Fig. S8). Further investigation might focus on whether this property helps prediction performance due to the varying contributions of the TCR and chain to specific binding across pMHCs [45].
@@ I.5 Supervised contrastive learning as a fine-tuning strategy
For this task we took all the TCRs annotated against the target pMHCs from our labelled TCR dataset, and split them into a training, a validation, and a testing set. We ensured that no study used for training or validation contributed any data to the test set, so that the fine-tuned model would not be able to achieve good performance simply by exploiting inter-dataset biases. The training set included 200 binders against each target pMHC, totalling to 1200 TCRs. The rest of the TCRs from the same studies were used to construct the validation set. TCRs from all remaining studies were used for the testing set, which comprised of 5670 TCRs. SCEPTR was fine-tuned on the training set with supervised contrastive learning, using the validation loss for early stopping (methods III.4).
We used the framework from section I.1 to benchmark the performance of fine-tuned SCEPTR, using the training set as the references. The results show that fine-tuning can greatly improve the ability of the model to discriminate between pMHCs (Fig. 5). Improvements are most noticeable for the pMHCs against which other methods achieve relatively low performance. When filtering all TCRs with greater than 90% or 80% sequence similarity to any training sequence from the test set, the fine-tuned model still improves performance significantly (Fig. S9) showing that learning goes beyond memorization of public TCRs.
Interestingly, unlike other models, fine-tuned SCEPTR makes better inferences by measuring the average distance between a query TCR and all reference TCRs instead only the nearest TCR (Fig. S13). This suggests an ability of supervised contrastive fine-tuning to help the model discover the commonalities between the multiple different binding solutions thought to exist for each pMHC. We thus analysed how fine-tuning changes the SCEPTR embedding distances between co-specific and cross-pMHC TCR pairs (Fig. 6). Unexpectedly, we found that a major difference of fine-tuned SCEPTR distances concerns cross-pMHC pairs. We observe that fine-tuning allows the model to identify a subset of “easy” negative pairs. These presumably involve TCRs the model is highly confident are specific to different pMHC, thus illustrating how discrimination between a fixed set of potential target pMHCs is easier than binary classification with respect to TCRs of arbitrary specificity. Conversely, the fine-tuned model’s performance degrades with respect to unseen pMHCs (Fig. S10), perhaps unsurprisingly given the very limited number of pMHCs represented in the training data.
Another notable feature of Fig. 5 is that performance varies substantially by ligand. That is, prediction is easier for some pMHCs, regardless of method. This is a phenomenon that has also been observed in public benchmarks [25]. Consequently, we used coincidence (order two Rényi) entropy measures [44, 45] to discover intrinsic properties of the pMHC-specific TCR repertoires that determine model AUROCs. We find that TCRs annotated against the “easier” pMHCs have lower V/J gene diversity (Fig. S11) and lower average sequence distances between pairs of TCRs (Fig. S12). This indicates that differences in models’ ability to predict TCR-pMHC specificity are linked to the diversity of epitope-specific repertoires.
@@ II Discussion
A limitation of our study is that we did not undertake a complete exploration of training and architectural hyperparameters, as well as training dataset choice. We envisage multiple avenues that may improve SCEPTR further. Firstly, training could be made more efficient by optimising the distribution of masked/dropped tokens during pre-training, taking into account the variable relevance of different parts of the sequence in determining specificity [45]. Secondly, as certain sequence motifs appear recurrently (e.g. CDR3 loops often begin with CAS), a more intelligent tokenisation scheme could offload learning of these primary sequence statistics into the tokenisation process. Finally, with the emergence of increasingly large paired-chain TCR datasets [57, 58, 59], retraining SCEPTR on data from multiple sources could eliminate biases inherent to specific experimental approaches and donor MHC restrictions.
@@
Uniformity incentivises the model to make use of the full representation space, while alignment minimises the expected distance between positive pairs (e.g. co-specific TCRs) [47]. From this view, contrastive learning on adaptive immune receptor data encourages PLMs to undo the large-scale distributional biases created by VDJ recombination through the uniformity term, while helping to identify features relating to TCR (co-)specificity via the alignment term. While autocontrastive learning approximates the alignment term through the generation of pairs of views, it still provides a direct empirical estimate for the uniformity term. Thus, a key benefit of autocontrastive learning may be that it reduces the confounding effects of VDJ recombination in embedding space. SCEPTR’s ability to “adjust” its distances for as demonstrated in Figs. 4 and S6 lends support to this conjecture.
@@
Another interesting avenue is the use of labels other than pMHC specificity – such as phenotypic annotations from single-cell data – as additional supervised contrastive training signals. While supervised contrastive learning does not currently lead to generalisable learning beyond training pMHCs, we expect a transition towards generalisation as larger volumes of specificity-labelled TCR data become available, as has been the case with supervised contrastive learning in other fields [67, 68, 69, 70, 52].
Finally, while the focus of this work given current data limitations has been on learning TCR embeddings, contrastive learning may also help us learn effective joint TCR-pMHC embeddings in the future when the joint space (and particularly the pMHC space) is better sampled, and thus ultimately enable the zero-shot prediction of TCR-pMHC specificity.
@@ III.1 Model benchmarking
For each pMHC, we varied the number of reference TCRs where . Within each model-pMHC--shot combination, we benchmarked multiple reference-test splits of the data to ensure robustness. For , we benchmarked every possible split. For , we benchmarked 100 random splits, where we ensured that the same splits were used across all models to reduce extraneous variance.
In assessing the statistical significance of differences in average model performance, we took a paired difference approach. We expected certain pMHCs and data splits to present a more difficult prediction problem than others. As we are interested in assessing the relative performance of models, we calculated the variance across splits of the difference between each individual model’s AUROC and the average across all models. For each model, we estimated this variance within each of the pMHCs, and then averaged these variances to obtain an estimate of overall variance.
@@ III.2 SCEPTR architecture
SCEPTR uses a simple, non-trainable embedder module, where a one-hot vector is used to encode token identity (22 dimensions for 20 amino acids plus special tokens <cls> and <mask>), and token positions are specified by first one-hot encoding the containing CDR loop number (6 dimensions), then encoding the token’s relative position within the loop as a single scalar variable (Fig. 2b). This results in initial token embeddings in , which are passed through a trainable linear projection onto . SCEPTR’s self-attention stack then operates at this fixed dimensionality (Fig. S16). SCEPTR’s self-attention stack comprises three layers, each with eight attention heads and a feed-forward dimensionality of 256, and is thus substantially simpler than existing models. Our tests suggest that relative position embedding helps SCEPTR learn better calibrated TCR co-specificity rules (see appendix C).
@@ III.3.1 Data
The unlabelled paired-chain TCR sequences used to pre-train SCEPTR were taken from a study by Tanno et al. [51], which provides 965,523 unique clonotypes sampled from the blood of 15 healthy human subjects. As opposed to traditional single-cell sequencing, Tanno et al. used a ligation-based sequencing method to resolve which chains paired with which chains. To mitigate potential noise from incorrect chain pairing, we applied an extra processing step to remove clonotypes that shared the same nucleotide sequence for either the or the chain, as previously described [44]. After filtering for functional TCRs using tidytcells, a TCR gene symbol standardiser [71], we retained 842,683 distinct clonotypes.